OpenStudy (anonymous):

Solve. (5x/7)+(3y/7)=1 x+(2y/3)=1 I keep getting fractions and I know that is not the right answer.

OpenStudy (andriod09):

is it dividing or is it fractions?

OpenStudy (anonymous):

factions

OpenStudy (andriod09):

\[(5x\div7)+(3y\div7)1\] or is \[(\frac{5x}{7})+(\frac{3y}{7})=1\]

OpenStudy (campbell_st):

use substitution by rewriting the 2nd equation as x = 1 - 2y/3 and substitute into equation 1 \[\frac{5(1 - \frac{2y}{3}}{7} +\frac{3y}{7} = 1\] multiply each term by 7 as the 1st step

OpenStudy (anonymous):

\[\frac{ 5x }{ 7 }+\frac{ 3y }{ 7 }=1\] \[x+\frac{ 2y }{ 3 }=1\]